OpenStudy (anonymous):

how do i show two 2x2 matrices are similar

OpenStudy (anonymous):

if the determinant of two matrices are same then they are said to be similar....

OpenStudy (amistre64):

if B = X' A X for some nonsingular matrix X, then A and B are said to be similar

OpenStudy (amistre64):

hmm, if we can develop a counter example to the det then not :) but i dont recall all the particulars

OpenStudy (amistre64):

similar matrixes have the same trace .. but that is not a sufficient condition

OpenStudy (anonymous):

oh ok ok i got it...

OpenStudy (anonymous):

So if I have two matrices, can I just find the eigenvalues or something to show that they are similar or do I have to actually find X?

OpenStudy (amistre64):

The tricky thing is this ... similar matrixes share alot, if not all, the same properties; but .... just because 2 matrixes share some particular properties does not guarantee us they are similar.

OpenStudy (amistre64):

B = X A X', or written another way, BX = XA B = ab X=wx A = mn cd yz pq BX = aw + by ax + bz cw + dy cx + dz XA = wm + xp wn + xq ym + zp yn + zq (a-m)w + by - px + 0z = 0 (a-q)x + bz - wn + 0y = 0 cw + (d-m)y - zp + 0x = 0 cx + (d-q)z - yn + 0w= 0 so, the matrix solution to X=wxyz would be the rref of: (a-m) -p b 0 -n (a-q) 0 b c 0 (d-m) p 0 c -n (d-q) or at least that is what i see it as ... which may mean that it isnt worth a hill of beans :)